# Revisiting Global Income Convergence in the 21st Century

### v1

@@ Abstract
This paper revisits the debate on income convergence between poor and rich countries. I challenge the view that there is little to no catch-up, and that changes in total factor productivity (TFP) drives cross-country income differences. Since 2000, income levels in poor countries have converged with rich countries at 0.8% annually, rising to 1.5% when excluding Sub-Saharan Africa. A growth accounting exercise incorporating capital income share heterogeneity shows that most convergence since 1980, and over half since 2000 outside Sub-Saharan Africa, results from convergence in physical and human capital inputs rather than TFP.
@@
Income Convergence, Capital Accumulation, Growth Accounting
@@ 1 Introduction
The growth literature traditionally finds little to no evidence of income levels in poor countries catching up to those in rich countries (unconditional convergence). It further attributes most income disparities across countries to total factor productivity (TFP) differences [12]. Moreover, growth accounting studies often assume a constant capital share of income of one-third across countries. This paper challenges these prevailing views. I find that after 2000, there was unconditional convergence at a rate of 0.8% for a broad sample of countries, and 1.5% when excluding Sub-Saharan Africa. This rate of unconditional convergence is in stark contrast to no unconditional convergence in [3], and higher than the convergence rate of 0.425% for the broad sample of countries and 1.175% outside Sub-Saharan Africa reported in [18]. More importantly, I find that convergence in physical capital and human capital, rather than TFP, has driven all the income convergence since 1980 and more than half of the convergence since 2000 outside Sub-Saharan Africa. Further, I find that capital income shares are not only higher than the commonly assumed one-third but also heterogeneous across countries, following an inverse U-shape across the income distribution.
I find different results compared to the previous literature primarily for two reasons. First, I include the time period post-2000 in my analysis where most of the convergence has actually taken place. In comparison, earlier growth accounting studies were based on a time period where divergence was the norm. Second, I incorporate heterogeneity of capital income shares as observed in the data. I show that assuming a constant capital income share across countries significantly reduces the role of physical capital and human capital in driving convergence, especially for the time period post-2000.
@@
The findings in the paper have important theoretical and policy implications. Economic growth models should incorporate heterogeneity in capital income shares and place greater emphasis on growth driven by the accumulation of capital. Since physical and human capital have been the primary drivers of income convergence, policymakers should prioritize investments in infrastructure, machinery, education, and healthcare, especially in Sub-Saharan Africa where growth has been slow.
This study builds on the papers by [18] and [15], who highlight the shift from divergence to convergence in recent decades. I add to the literature by examining the proximate correlates correlates of convergence within a traditional growth accounting framework, focusing on the roles of physical capital, human capital, and TFP. Earlier growth accounting literature, such as [14], [7], and [12], focused on periods when divergence between countries was the norm, attributing most growth to TFP. By incorporating heterogeneity in capital shares and focusing on a period marked by convergence, I provide new insights into the factors driving income convergence, highlighting the significant roles of physical and human capital accumulation.
@@ 2 Data
adjusted for purchasing power parity (PPP) ($rgdpo$), population ($pop$), human capital ($hc$), and labor share of income ($labsh$) for all countries come from the Penn World Tables (PWT10.01). [^2] See [8] for details. Regional classifications are based on the World Bank’s classification of economies. I calculated the capital-output ratio as the ratio of real capital stock (rnna) to real GDP (rgdpna). I use per capita GDP, adjusted to 2017 PPP dollars, as the measure of income ($rgdpo/pop$).
@@
The main sample consists of a balanced panel of 131 countries from 1970 to 2019. The main sample excludes countries with small populations and those whose GDP is heavily dependent on oil rents, which is standard practice in the literature. [^3] Specifically, I exclude countries whose population never exceeded 0.2 million at any point in time. I also exclude countries where oil rents accounted for more than 50% of GDP at any time. The capital measure in the PWT includes fixed reproducible capital but excludes natural resource capital, such as land, inventories, subsoil assets, or intangible capital. This difference is particularly significant for oil-rich countries, which is why they are excluded from the main analysis. [15] also excludes small countries, and oil rich countries in their analysis. I use a sample of 84 economies outside Sub-Saharan Africa, for which the PWT reports human capital, physical capital, and capital share of income to conduct the main growth decomposition exercise. [^4] For variance measures, I further exclude Venezuela. Venezuela is excluded for statistical reasons, as it experienced an economic crisis beginning in 2000, with current GDP per capita lower than its 1960 level. This makes Venezuela an outlier with a large influence on variance measures. The time series of Venezuela’s GDP per capita is plotted in the supplementary appendix for reference.
3 Rate of Income Convergence
This section provides evidence that, post-2000, initially poor countries have been growing faster than initially rich countries. This phenomenon, where poorer economies tend to grow more rapidly than wealthier ones, is commonly referred to as unconditional convergence (or $\beta$-convergence) in the growth literature. I quantify the rate at which poorer countries have been catching up with their richer counterparts. Formally, I calculate the rate of $\beta$-convergence between time period $t$ and $T$ by estimating the following regression:
$\frac{1}{T}(\log(y_{i,T})-\log(y_{i,t}))=\gamma+\beta\log(y_{i,1980})+\epsilon_{i,t}.$ (1)
@@ 3 Rate of Income Convergence
Here, $y_{i,t}$ refers to the income of country $i$ in year $t$, and $\epsilon_{i,t}$ captures the shocks to the annualized growth rate between time periods $t$ and $T$. I assume that these shocks are independently and identically distributed (iid) with a normal distribution. I estimate separate regressions for the periods 1980-2000 and 2000-2019. In this specification, $\beta$ is identified using cross-sectional variation in annual growth rates across countries initially at different income levels. Additionally, note that the regression in equation 1 includes no country-level controls, meaning that the $\beta$ identified here reflects the rate of unconditional convergence. In both cases, I fix the base-year income at its 1980 level. Supplementary Appendix provides additional analysis, running the convergence regression with 1960 as the base year.
@@
Table 1 presents the estimates for the rate of $\beta-$convergence. Columns (1) and (2) report the estimates for the full balanced panel of countries, while Columns (3) and (4) show results for the set of countries excluding Sub-Saharan Africa. In Column (1), the convergence coefficient is positive and significant for the full sample of countries during the 1980-2000 period. This indicates that rich countries were initially growing faster than poor countries, leading to increased income divergence. However, for countries outside Sub-Saharan Africa, the convergence coefficient is statistically indistinguishable from zero during the same period, providing no evidence of convergence in either sample.
@@
In contrast, the rate of convergence has seen a remarkable reversal in the past two decades. In the full sample, initially poor countries have been catching up to initially rich countries at an annual rate of 0.8%. For countries outside Sub-Saharan Africa, the catch-up rate has been even faster, at 1.5% annually. For comparison, [18] found a the rate of convergence of 0.425% for the full sample of countries and 1.175% outside Sub-Saharan Africa between 2000-2019 with 2000 as the base year. [15] don’t report the rate of convergence outside Sub-Saharan Africa but do find the rate of convergence to be 0.7% in broad sample of countries between 2005-2015 with 2005 as the base year. [1] estimates the rate of conditional convergence to be 1.7% regressing 5-year growth rates on lagged income level and country level-time varying controls. The unconditional convergence rate of 1.5% observed here is substantial in comparison. This rate of $\beta-$convergence implies a half-life of 46 years, meaning that, starting from 2019, poor countries outside Sub-Saharan Africa would close half the income gap with developed countries by 2065, assuming all countries share the same steady-state income level.  [^5] The half-life is calculated as $\tau=-\ln(2)/\ln(1-\beta)$. Moreover, nearly 24% of the variation in growth rates across countries outside Sub-Saharan Africa can be explained by their income levels in 1980. Macro growth regressions typically suffer from low explanatory power since there are a lot of variables determining future growth trajectories not included in the model. The relatively high $R^{2}$ achieved here without the inclusion of additional covariates, suggests that a country’s position in the income distribution is a strong proxy for a range of other factors affecting subsequent growth rates.
Given that most of the convergence post-2000 is observed outside Sub-Saharan Africa, and that data for countries outside Sub-Saharan Africa is generally more reliable [21], the remainder of the paper focuses on this subset of countries. Results for the full set of countries are provided in the Supplementary Appendix.
4 Trends in cross country income differences
This section documents that the decline in overall dispersion of income among counties outside Sub-Saharan Africa started to decline post 2000. Table 2 presents the evolution of income dispersion across countries outside Sub-Saharan Africa over time, using various measures. Columns 1, 2, and 3 display the P90/P10, P90/P50, and P50/P10 percentile ratios of GDP per capita, which capture different aspects of income inequality across the distribution. Column 4 reports the variance of the log of GDP per capita, and Column 5 shows the income ratio between the top 5 and bottom 5 countries.
The data reveal that all measures of income dispersion increased between 1980 and 1990. While the P90/P10 and P50/P10 ratios began to decline after 1990, the most notable trend is the consistent decrease in all measures of income inequality starting from 2000. By 2019, each measure of income dispersion reached its lowest point in the last 40 years.
@@ 4 Trends in cross country income differences
Given the significant decline in cross-country income dispersion, the next section conducts an accounting exercise to decompose this convergence into parts attributable to convergence in TFP, the capital-output ratio, and human capital.
@@ 5 Accounting for Decline in Income Dispersion
In this section, I perform an accounting exercise to attribute the decline in income dispersion to reductions in TFP differences and input dispersion. Unlike many standard approaches in the literature, which assume a uniform capital share parameter of one-third, this accounting exercise based on percentile ratios allows the capital share of income to vary continuously with the income percentile, as observed in PWT10.01. The accounting exercise using variance as a measure of income dispersion is presented in the supplementary appendix.
Let $p$ denote the income percentile. Suppose that per capita output ($y$) is a function of physical capital ($k$), human capital ($h$), and a residual term ($A$). [^6] It is common to interpret the residual term ($A$) as total factor productivity (TFP), but more broadly, it includes both observable and unobservable factors that affect output beyond the inputs. The output production function, parameterized by $p$, takes the following form:
$y(p)$ $=A(p)(k(p))^{\alpha(p)}(h(p))^{1-\alpha(p)}$
This is the standard Cobb-Douglas production function, with the modification that the capital share ($\alpha$) is allowed to vary across the income distribution. This flexibility is motivated by the large variation in capital share across countries and over time. Figure 1 shows the capital share along the income distribution, as computed from PWT10.01. The capital share is not only heterogeneous but higher than what is typically assumed in the literature. The production function can be rearranged and written in intensive form as follows:
$y(p)=A(p)^{\frac{1}{1-\alpha(p)}}(\frac{k}{y}(p))^{\frac{\alpha(p)}{1-\alpha(p)}}(h(p)).$
Writing the production function in this form reduces the effect of differences in per capita capital ($k$), since per capita capital is allowed to adjust endogenously in response to changes in total factor productivity ($A$). [^7] As argued in [14], under firm optimization, equalization of the marginal product of capital with the rental rate means that an increase in total factor productivity (TFP) leads to a mechanical increase in capital, assuming a constant rental rate. Since my goal is to separately identify the contributions of total factor productivity (TFP) growth and capital accumulation to output growth, I prefer to use the production function written in this form. Given this functional form, the income difference between percentiles $p_{0}$ and $p_{n}$ can be decomposed as follows:

### v2

@@ Abstract
Recent research has documented a reversal from divergence to convergence in income levels between rich and poor countries after 2000. This paper employs a growth accounting framework to investigate the proximate sources of convergence over the period 1980-2019. I find that while output levels began to converge only after 2000, capital (physical and human) had already been converging during 1980-2000. The divergence in total factor productivity (TFP) during this earlier period offset the gains from capital convergence, resulting in little overall income convergence. After 2000, capital maintained its pattern of convergence, and, unlike in the earlier period, TFP also began to converge. Quantitatively, more than half of the convergence between 2000 and 2019—and nearly all the convergence between 1980 and 2019 outside Sub-Saharan Africa—can be attributed to the convergence in capital rather than to improvements in TFP. These findings highlight that factor accumulation has played a much larger role in driving long-run convergence dynamics than previously documented in the literature.
@@
Income Convergence, Capital Accumulation, Growth Accounting, Total Factor Productivity
@@ 1 Introduction
The traditional literature on economic growth has long debated whether poorer countries’ incomes converge toward those of richer countries. Empirical studies, particularly those prior to the early 2000s, generally found limited or no support for unconditional convergence [3, 18]. However, recent studies by [17] and [14] challenge this view, documenting a reversal from divergence to convergence in recent decades. This documented reversal prompts two critical yet relatively unexplored questions: first, what factors can explain this turnaround in convergence dynamics; and second, which proximate determinants have contributed significantly to the observed income convergence?
I address these questions by first replicating the empirical finding of convergence in recent decades, thereby establishing a consistent empirical basis for further analysis. I then examine the proximate drivers of this shift using a growth accounting framework that decomposes income growth into contributions from total factor productivity (TFP), physical capital, and human capital. In doing so, I explicitly account for heterogeneity in capital income shares across countries, a feature that standard growth accounting exercises often abstract from. This approach allows for a more accurate quantification of the role played by each factor in driving convergence, and highlights how their relative contributions have changed over time.
I find that the reversal from divergence to convergence can be attributed to two main factors: first, total factor productivity (TFP) shifted from a diverging pattern before 2000 to convergence thereafter; and second, physical and human capital inputs continued to converge consistently. While TFP transitioned from divergence to convergence post-2000, its contribution to long-run income convergence remains secondary. I show that over the period 2000-2019, more than half of the observed income convergence among countries outside Sub-Saharan Africa (SSA) arises from convergence in physical and human capital. More strikingly, over the longer timeframe of 1980-2019, income convergence among countries outside SSA occurs despite divergence in total factor productivity (TFP). The entirety of the long-run convergence in output levels for these countries is driven solely by narrowing gaps in physical and human capital accumulation. Additionally, I find that for poorer countries within Sub-Saharan Africa (SSA), there has been no convergence in the capital-output ratio since 2000. Notably, while the extent of total factor productivity (TFP) convergence between rich economies and poor countries outside SSA mirrors that observed between rich economies and poor countries in SSA, the markedly lower degree of output convergence in the latter is entirely accounted for by the absence of capital convergence. This pattern suggests that the lack of capital convergence is an important factor contributing to the region’s relative economic stagnation.
These findings provide a revised understanding of the drivers behind global income dynamics. Prior growth accounting and development accounting exercises consistently highlighted differences in total factor productivity (TFP) as the central factor explaining persistent income gaps and divergence across countries [13, 6, 11]. While TFP indeed remains the dominant explanatory factor in the divergence era preceding 2000, its relative importance diminishes significantly thereafter. Specifically, for the period following 2000, and even more clearly over the longer span from 1980 to 2019, the observed convergence in global income levels is primarily driven by convergence in physical and human capital accumulation rather than TFP. The differences between these findings and the prior literature primarily reflect two key factors: first, this paper explicitly analyzes the more recent period of convergence post-2000, whereas earlier studies predominantly considered periods of divergence; second, the analysis incorporates heterogeneity in capital income shares across countries and over time. Assuming a constant capital income share underestimates the contribution of factor inputs, especially in the period following 2000. This paper also contributes to the literature on economic growth in Sub-Saharan Africa (SSA) by identifying the non-convergence of capital accumulation as a critical factor underlying the region’s persistent income stagnation.
@@
The findings in this paper have important theoretical and policy implications. Economic growth models should incorporate heterogeneity in capital income shares and place greater emphasis on growth driven by the accumulation of physical and human capital. Given that these factors have been the primary drivers of income convergence, policymakers should prioritize investments in infrastructure, machinery, education, and healthcare—especially in regions such as Sub-Saharan Africa where growth has been slow.
@@ 2 Data
This section summarizes the data used in the current study. Data for real GDP adjusted for purchasing power parity (PPP) ($rgdpo$), population ($pop$), human capital ($hc$), and labor share of income ($labsh$) for all countries come from the Penn World Tables (PWT10.01). [^1] See [7] for details. Regional classifications are based on the World Bank’s classification of economies. I calculate the capital-output ratio as the ratio of real capital stock (rnna) to real GDP (rgdpna). I use per capita GDP, adjusted to 2017 PPP dollars, as the measure of income ($rgdpo/pop$). The selection of output side real gdp (rgdpo) over alternative Penn World Table output measures follows the recommendations of [7] for comparing productive capacities across countries and over time.
@@
All samples excludes countries with small populations and those whose GDP is heavily dependent on oil rents, which is standard practice in the literature. [^2] Specifically, I exclude countries whose population never exceeded 0.2 million at any point in time. I also exclude countries where oil rents accounted for more than 50% of GDP at any time. The capital measure in the PWT includes fixed reproducible capital but excludes natural resource capital, such as land, inventories, subsoil assets, or intangible capital. This difference is particularly significant for oil-rich countries, which is why they are excluded from the main analysis. [14] also excludes small countries, and oil rich countries in their analysis. I use a sample of 84 economies outside Sub-Saharan Africa, for which the PWT reports human capital, physical capital, and capital share of income to conduct the main growth decomposition exercise. [^3] For variance measures, I further exclude Venezuela. Venezuela is excluded for statistical reasons, as it experienced an economic crisis beginning in 2000, with current GDP per capita lower than its 1960 level. This makes Venezuela an outlier with a large influence on variance measures. The time series of Venezuela’s GDP per capita is plotted in the supplementary appendix for reference.
3 Income Convergence
I replicate the empirical framework of [17] and [14] to investigate whether, in the post-2000 period, economies with lower initial income per capita have tended to grow more rapidly than those with higher initial incomes. This pattern-commonly referred to as unconditional (or $\beta$-) convergence-implies that poorer countries catch up to richer ones over time. To quantify the speed of convergence between dates $t$ and $T$, I estimate the following specification:
$\frac{1}{s}\ln\Big{(}\frac{y_{i,t+s}}{y_{i,t}}\Big{)}=\beta_{0}-\Big{(}\frac{1-\exp^{\beta s}}{s}\Big{)}\ln(y_{i,t})+\epsilon_{i,t+s}.$ (1)
@@ 3 Income Convergence
Here, $y_{i,t}$ refers to the income of country $i$ in year $t$, $y_{i,t+s}$ refers to the income of country $i$ in year $t+s$, and $\epsilon_{i,t+s}$ captures the shocks to the annualized growth rate between time periods $t$ and $t+s$. I assume that these shocks are independently and identically distributed (iid) with a normal distribution. I estimatate specification 1 using non-linear least square with heteroskedaticity-robust standard errors. I estimate separate regressions for the periods 1980-2000 and 2000-2019. In this specification, $\beta$ is identified using cross-sectional variation in annual growth rates across countries initially at different income levels. Additionally, note that the regression in equation 1 includes no country-level controls, meaning that the $\beta$ identified here reflects the rate of unconditional convergence.
@@
Table 1 presents estimates of the rate of $\beta$-convergence. Columns (1) and (2) report results for the full balanced panel, whereas columns (3) and (4) restrict the sample by excluding Sub-Saharan African economies. Over the 1980-2000 period, the coefficient in column (1) is positive and statistically significant, indicating that wealthier countries grew faster than poorer ones and thereby increasing income divergence. The coefficient in column (3) does not differ significantly from zero, showing that even when Sub-Saharan Africa is excluded, there was no convergence over the 1980-2000 period.
In contrast, the rate of convergence has seen a remarkable reversal in the past two decades. In the full balanced panel, initially poorer economies have closed their income gap at an annual rate of 0.62 percent. When Sub-Saharan African countries are excluded, the catch-up rate rises to 1.5 percent per year. By comparison, [17] report a 0.425 percent convergence rate for 2000-2019 in their full sample and 1.35 percent outside Sub-Saharan Africa, while [14] document a 0.7 percent rate for 2005–2015 (base year 2005) in a broad country panel. The somewhat higher coefficient estimated here reflects differences in sample composition and output measurement: [17] employ expenditure-side real GDP (rgdpe) and omit countries with populations below one million, whereas this analysis uses output-side real GDP (rgdpo) and excludes economies with fewer than 0.2 million inhabitants, in line with [14]. Table 3 presents the convergence coefficients across alternative Penn World Table output series for both samples.
@@
The unconditional convergence coefficient for the post 2000 period outside SSA is also close to the rate of conditional convergence found in recent literature. [1] estimates the rate of conditional convergence to be 1.7% regressing 5-year growth rates on lagged income level and country level-time varying controls. The unconditional convergence rate of 1.5% observed here for countries outside SSA is substantial in comparison. This rate of $\beta-$convergence rate of 1.5% implies a half-life of 53 years, meaning that, starting from 2020, poor countries outside Sub-Saharan Africa would close half the income gap with developed countries by 2073, assuming all countries share the same steady-state income level.  [^4] The half-life is calculated as $\tau=-\ln(2)/\ln\Big{(}1-\frac{1-\exp^{\beta s}}{s}\Big{)}$.
@@
All Countries
1980 18.56 5.30 3.50 1.15 18.49
1990 24.95 5.55 4.50 1.37 23.26
2000 31.95 6.88 4.64 1.62 29.53
2010 27.63 4.51 6.12 1.54 30.71
2019 25.06 4.41 5.69 1.47 27.17
Beta-convergence is necessary but not sufficient for sigma-convergence [22, 3]. [^5] Sigma convergence refers to a decline over time in the cross-sectional dispersion of per-capita income across economies. Consequently, even if initially poorer economies grow more rapidly than their richer counterparts, the cross-sectional dispersion of per-capita income need not decline. Table 2 reports five measures of income dispersion-namely the P90/P10, P90/P50 and P50/P10 percentile ratios of GDP per capita; the variance of log GDP per capita; and the ratio of the top five to bottom five incomes-for each decade from 1980 to 2019. The table enables a clear assessment of whether-and when-sigma-convergence accompanied the episodes of unconditional convergence identified earlier.
For countries outside SSA, income dispersion widened during the 1980s—by 1990 the P90/P10 ratio had risen to nearly 20 and the variance of log GDP per capita to above 1.1. Since 2000, every measure has fallen steadily: by 2019 the P90/P10 ratio had roughly halved to below 10, and the variance of log GDP per capita declined to 0.8, reaching its lowest level in four decades. Hence, after 2000, initially poorer non-SSA economies not only grew faster than their richer counterparts but also saw a marked reduction in overall income dispersion—demonstrating both beta- and sigma-convergence in that subsample.
For the full country sample, dispersion rose sharply through the 1980s and 1990s, with the P90/P10 ratio climbing from 18.6 in 1980 to nearly 32.0 in 2000 before easing to 25.1 by 2019. Likewise, the variance of log GDP per capita peaked at 1.62 in 2000 and then fell modestly to 1.47 by 2019. By contrast, the P50/P10 ratio rose steadily from 3.50 in 1980 to 5.69 in 2019, signaling a widening gap between middle-income and poorest countries. These trends imply only limited sigma-convergence: although top-to-bottom dispersion declined somewhat after 2000, overall income inequality in 2019 remains well above its 1980 level.
Table 2 demonstrates that, although economies outside Sub-Saharan African achieved both beta- and sigma-convergence after 2000, many Sub-Saharan African countries have continued to stagnate. Since most of the observed convergence-and the most reliable data-resides outside SSA [21], the subsequent analysis focuses on non-SSA economies. Full-sample results appear in the Appendix.
@@ 4 Accounting for Decline in Income Dispersion
In this section, I perform an accounting exercise to attribute the decline in income dispersion to reductions in TFP differences and input dispersion. Unlike many standard approaches in the literature, which assume a uniform capital share parameter of one-third, this accounting exercise based on percentile ratios allows the capital share of income to vary continuously with the income percentile, as observed in PWT10.01. The accounting exercise using variance as a measure of income dispersion is presented in the appendix.
Let $p$ denote the income percentile. Suppose per-worker output is generated by a constant-returns-to-scale technology
$y(p)=F\bigl{(}A(p),k(p),h(p)\bigr{)},$
where $A(p)$ is Hicks-neutral productivity. [^6] It is common to interpret $A(p)$ as total factor productivity (TFP), but more broadly it includes both observable and unobservable factors that affect output beyond the inputs. As argued in [13], under firm optimization, equalization of the marginal product of capital with the rental rate implies that TFP growth mechanically raises the capital stock under a constant rental rate; to separate TFP and capital-accumulation effects in output growth, the production function is written in intensive form. Then the difference in log-output between percentiles $p_{0}$ and $p_{n}$ admits the decomposition